OpenStudy (loser66):

Without using chi-square, if \(M(t)= e^{500t +5000t^2}\) find \(P(27060\leq (X-500)^2\leq 50240)\) Please, help

OpenStudy (loser66):

@kirbykirby

OpenStudy (loser66):

From moment generating function, we have \(\mu_x =500\) and \(\delta ^2=10000\) then??

OpenStudy (perl):

you cant integrate the function ?

OpenStudy (loser66):

page113 , problem 3.3-12

OpenStudy (zarkon):

solve the inequality for X the mgf is from a normal distribution

OpenStudy (loser66):

but X is not just X, it is (X-500)^2, if applying \(Z=\dfrac{X-\mu}{\delta}\), is it not that using chi-sqruare??

OpenStudy (zarkon):

solve the inequality for X...then standardize it

OpenStudy (zarkon):

and it is \(\sigma\) not \(\delta\)

OpenStudy (loser66):

:)

OpenStudy (zarkon):

you are using the chi square if you work with \(Z^2\) you are just going to work with Z

OpenStudy (loser66):

\((P({2.7060}\leq Z^2\leq 5.0240\\(P(\sqrt{2.7060}\leq Z\leq \sqrt{5.0240}\) for N (0,1) Am I right?

OpenStudy (zarkon):

Z could be negative for example if \[1<z^2<4\] then \(1<z<2\) and \(-2<z<-1\)

OpenStudy (zarkon):

not 'and' ... it's 'or'

OpenStudy (loser66):

But they have the same probability, right?

OpenStudy (loser66):

the same mean and variance also.

OpenStudy (zarkon):

yes...same prob...so you have to multiply by two

OpenStudy (loser66):

I got it. Thank you. :)